# VerifiableFL: Verifiable Claims for Federated Learning using Exclaves

### v3

@@ 7.3. Scalability
First, we measure VerifiableFL’s scalability with the GPT2 model for language modeling (Fig. 8(a)). Here, we scale VerifiableFL from 1 to 4 clients, each with a GPU. With $3\times$ increase in FL clients, VerifiableFL, NVFlare (CVM) and NVFlare (No security) have a $\times, 2.8×2.8\times$, and $2.9\times$ performance increase, respectively.
We also consider the BERT model for named-entity recognition (Fig. 8(b)), and the VGG9 model for image classification (Fig. 8(c)). In these experiments, we increase the number of CPUs when scaling clients. With the BERT model, a $3\times$ increase in FL clients leads to a $2.8\times$, $3.0\times$, and $2.9\times$ increase in throughput for VerifiableFL, NVFlare (CVM), and NVFlare (No security), respectively. For the VGG9 model, we observe that the throughput for VerifiableFL, NVFlare (CVM), and NVFlare (No security) increases by $2.9\times$, $3.0\times$, and $2.8\times$, respectively, when we quadruple the client count.
Both results show that VerifiableFL scales effectively with more FL clients. Since all clients run in parallel, the runtime and I/O overheads of clients do not accumulate.
Model Component Time (s)
GPT2 Training Exclave 8,659 73%
DP Exclave 1,176 10%
Aggregation Exclave 50 ¡1%
Model Update Exclave 29 ¡1%
FL Framework 1,982 17%
VGG9 Training Exclave 4,642 97%
DP Exclave 21 ¡1%
Aggregation Exclave 2 ¡1%
Model Update Exclave 1 ¡1%
FL Framework 124 3%
@@ 7.4. Performance breakdown
We also break down the execution time of VerifiableFL’s components when running an FL job from §7.2. Tab. 2 shows the performance breakdown. With the GPT2 and VGG9 training tasks, training exclave dominates execution time (97% for VGG9; 73% for GPT2). Note that GPT2 trains with GPU exclaves, while VGG9 trains with CPU exclaves. Thus, the GPT2 training task occupies a smaller proportion of the overall time.
@@
The time spent on executing the aggregation and model update exclaves increases linearly with the model size, because the averaging/addition operations they perform have linear time complexity. Note that the DP exclave has a worst-case quadratic time complexity, because the SVT-DP algorithm iterates over all model parameters potentially multiple times in order to calculate the noise and exclude parameters beyond the noise threshold in the iteration. The total time spent on the DP, aggregation, and model update exclaves is less than 1% for VGG9 and up to 10% for GPT2.
@@
The execution time for the FL framework, including the time spent on orchestration and its communication, also increases linearly with the model size. Frequent communication and model (de)serialization makes the FL framework the second largest contributor to VerifiableFL’s execution time: 17% and 3% for GPT2 and VGG9, respectively. This overhead comes from VerifiableFL’s orchestration of tasks that execute in exclaves: it is necessary for VerifiableFL to deploy exclaves in separate CVMs to achieve the fine-grained attestation of each exclave needed for its transparency guarantees.
7.5. Data sanitization overhead
@@
We evaluate the overhead of VerifiableFL’s data sanitization exclave, which a data provider can deploy and run prior to FL training. The data sanitization exclave is deployed as an Azure NCC40ads_H100_v5 CVM. It runs a harmful-text classification model, granite-guardian-hap-38m (Padhi et al., 2024), to detect toxic and abusive content in the dataset; the exclave then removes the identified harmful content to sanitize the potentially contaminated dataset.
@@
We evaluate the time cost of data sanitization exclave when sanitizing each data provider’s Wikitext103 dataset. The result reveals that the data sanitization exclave takes 257 seconds to sanitize each data owner’s dataset, adding 2% overhead to VerifiableFL’s GPT2 training.
7.6. Runtime attestation overhead
@@
In this micro-benchmark, we measure the time overhead of runtime attestation. We use a single CVM with 4 vCPUs and 16 GB of memory. We then generate a SHA256 digest over 1 GB of input data and include the resulting hash digest as the runtime claim of the attestation request.
@@
On average, creating and signing an attestation report with vTPM takes 3 ms. This step has constant time complexity, because both creation and signing operations take a fixed-size hash digest as input. Generating hash digests is the most time-consuming operation during attestation: it takes 686 ms to calculate a SHA-256 hash over the 1 GB input data, and the time grows linearly with the input data size.
@@
We also measure the attestation overhead with PSP: the CVM requires less than 7 ms to fetch a signed attestation report from the PSP. Although we use vTPM attestation to emulate PSP attestation, their performance difference is minimal (under 4 ms per attestation) and negligible compared to VerifiableFL’s end-to-end training, which takes hours.
@@
7.7. Secure storage overhead
@@
We also explore the performance of VerifiableFL’s secure storage. We mount a disk protected by dm-verity on an Azure DC32as_v5 CVM (P30 data disk with up to 200 MB/s and 5,000 IOPS). We then use the FIO benchmark (Axboe, [n. d.]), which sequentially reads data from the ext-4 file system on the secure disk. We compare against insecure storage as a baseline, which uses a standard D32as_v5 VM on an unprotected Azure P30 disk.
@@
As we shown in Fig. 9, the sequential read throughput of the integrity-protected disk is 62% lower compared to regular storage. In terms of read latency, the secure storage module adds $1.7\times$ extra latency when reading a block from disk to memory. This is due to the cost of checking the block against the Merkle tree. The resulting read amplification impacts both throughput and latency, which adds an end-to-end performance cost to VerifiableFL and NVFlare (CVM).
@@ 8. Conclusions
We described VerifiableFL, a new system for executing FL jobs, while enabling an auditor to verify claims on the final trained FL model. VerifiableFL introduces exclaves as a new abstraction for the execution of FL tasks with integrity, which allows it to attest all performed data transformations at runtime. Since all executed FL tasks are attested through exclave data records (EDRs), it becomes impossible for participants to deviate from the correct FL training job, thus violating model claims, without detection. Exclaves produce EDRs, which can be combined to form an exclave dataflow graph (EDG) that spans the entire FL job. An auditor uses the EDG to verify end-to-end claims of the FL job. We showed that VerifiableFL can be integrated with a state-of-the-art FL framework with modest code changes, while detecting a wide range of deviations with low overhead.

### v4

@@ 7.3. Scalability
First, we measure VerifiableFL’s scalability with the GPT-2 model for language modeling (Fig. 8(a)). Here, we scale VerifiableFL from 1 to 4 clients, each with a GPU. With $3\times$ increase in FL clients, VerifiableFL, NVFlare (CVM) and NVFlare (No security) have a $\times, 2.8×2.8\times$, and $2.9\times$ performance increase, respectively.
We also consider the BERT model for named-entity recognition (Fig. 8(b)), and the VGG9 model for image classification (Fig. 8(c)). In these experiments, we increase the number of CPUs when scaling the number of clients. With BERT, a $3\times$ increase in FL clients leads to a $2.8\times$, $3.0\times$, and $3.0\times$ increase in throughput for VerifiableFL, NVFlare (CVM), and NVFlare (No security), respectively. For VGG9, we observe that the throughput for VerifiableFL, NVFlare (CVM), and NVFlare (No security) increases by $2.9\times$, $3.0\times$, and $3.0\times$, respectively, when we quadruple the client count. Both results show that VerifiableFL scales effectively with more FL clients, similar to other baselines. Since all clients run in parallel, the runtime and I/O overheads of clients do not accumulate.
@@
Component GPT2 VGG9
time (sec) time (sec)
Training Exclave 8,659 (73%) 4,641 (97%)
DP Exclave 1,176 (10%) 21 (¡1%)
Aggregation Exclave 50 (¡1%) 1 (¡1%)
Model Update Exclave 29 (¡1%) 1 (¡1%)
FL Framework 1,983 (17%) 124 (3%)
@@ 7.4. Performance breakdown
We also provide a breakdown the execution time of VerifiableFL’s components when running the FL job from §7.2. Tab. 2 shows the results. With the GPT-2 and VGG9 training tasks, the training exclave dominates execution time (97% for VGG9; 73% for GPT-2). Note that GPT-2 trains with GPU exclaves, while VGG9 trains with CPU exclaves. Thus, the GPT2 training task occupies a smaller proportion of the overall time.
@@
The time spent on executing the aggregation and model update exclaves increases linearly with the model size, because the performed averaging/addition operations have linear time complexity. Note that the DP exclave has a worst-case quadratic time complexity, because the SVT-DP algorithm iterates over all model parameters (potentially multiple times) to calculate the noise and exclude parameters beyond the noise threshold in the iteration. The total time spent on the DP, aggregation, and model update exclaves is less than 1% for VGG9 and up to 11% for GPT-2.
@@
The execution time for the FL framework, including the time spent on orchestration and communication, also increases linearly with model size. Frequent communication and model (de)serialization makes the FL framework the second largest contributor to VerifiableFL’s execution time: 17% and 3% for GPT-2 and VGG9, respectively. This overhead comes from VerifiableFL’s orchestration of tasks that execute in exclaves: it is necessary for VerifiableFL to deploy exclaves in separate CVMs to provide the fine-grained runtime attestation of each exclave done by the hardware.
@@
In addition, we evaluate the overhead of VerifiableFL’s data sanitization exclave, which a data provider deploys and runs prior to FL training on an Azure NCC40ads_H100_v5 CVM. The data sanitization exclave uses a classification model for harmful text, granite-guardian-hap-38m (granite), to detect toxic or abusive content and then removes the identified content to sanitize the dataset.
@@
We measure the execution time of the data sanitization exclave when sanitizing each data provider’s Wikitext103 dataset. The result shows that the data sanitization exclave takes 257 seconds to sanitize each data owner’s dataset. This adds 2% overhead to VerifiableFL’s GPT-2 training. We conclude that enforce datasets properties using standard sanitisation techniques therefore is feasible in practice.
7.5. Overhead of runtime attestation
@@
In this micro-benchmark, we measure the overhead due to runtime attestation. We use a single CVM with 4 vCPUs and 16 GB of memory to generate a SHA256 digest over 1 GB of input data and include the resulting hash digest as the runtime claim in an attestation request.
@@
We observe that, on average, creating and signing an attestation report with vTPM takes 4 ms. This step has constant time complexity, because both creation and signing operations take a fixed-size hash digest as input. Generating hash digests is the most time-consuming operation during attestation: it takes 687 ms to calculate a SHA-256 hash over the 1 GB input data, and the time grows linearly with the input data size.
@@
We also measure the attestation overhead with AMD’s PSP: the CVM requires less than 7 ms to fetch a signed attestation report from the PSP. Although we use vTPM attestation to emulate PSP attestation, the performance difference between them is minimal (under 3 ms per attestation) and negligible compared to VerifiableFL’s end-to-end training time, which takes hours.
@@
7.6. Overhead of secure storage
@@
We also explore the performance impact of VerifiableFL’s secure storage for datasets. We mount a disk protected by dm-verity on an Azure DC32as_v5 CVM (P30 data disk with up to 200 MB/s and 5,000 IOPS). We then use the FIO benchmark (fio), which sequentially reads data from the ext-4 file system on the secure disk. We compare against insecure storage as a baseline, which uses a standard D32as_v5 VM on an unprotected Azure P30 disk.
@@
As Fig. 9 shows, the sequential read throughput of the integrity-protected disk is 62% lower compared to regular storage. This is due to the cost of checking the block against the Merkle hash tree. Latency also increases accordingly: the secure storage module adds an additional $1.7\times$ overhead over the baseline latency (reported latencies include queuing time from the in-flight I/Os). The resulting read amplification impacts both throughput and latency, which adds an end-to-end performance cost to VerifiableFL.
Table 3. Overhead of auditing
#clients $\times$ Signature EDG total storage
#rounds verification (ms) creation (ms) (ms) (MB)
10 $\times$ 10 26 3 29 ¡1
10 $\times$ 100 196 13 209 2
10 $\times$ 1,000 1,886 131 2,017 17
100 $\times$ 10 218 20 238 2
100 $\times$ 100 1,802 146 1,948 16
100 $\times$ 1,000 17,404 1,837 19,241 162
7.7. Overhead of auditing
We evaluate the overhead that the auditing process incurs in real-world settings, when the number of clients scales from 10 to 100, and the number of rounds scales from 10 to 1,000. We generate synthetic EDRs and measure the auditor’s computation and storage overheads in a standard VM with 4 vCPUs and 16 GB memory.
Tab. 3 shows the results. When 100 clients train for 1000 rounds with VerifiableFL, it takes the auditor 19 seconds to verify all EDR signatures and construct the EDG, which is negligible in comparison to the FL training time. The incurred storage cost during auditing is 162 MB, which is also small. We conclude that the auditing process of VerifiableFL only introduces a small overhead, even with larger number of clients and training rounds.
8. Related Work
Mitigating attacks in FL. Various aggregation protocols are designed to mitigate privacy and integrity attacks in FL. Secure aggregation (bonawitz2017secure; sharma2019secure) addresses FL privacy concerns when aggregators are untrusted through cryptographic means. For example, Bonawitz et al. (bonawitz2017secure) implement secure aggregation using pair-wise Diffie-Hellman key agreement and t-out-of-n Shamir’s secret sharing, whilst Sharma et al. (sharma2019secure) use additive secret sharing. These protocols are complementary to VerifiableFL because they ensure the privacy of the FL training process, while VerifiableFL provides verifiable integrity guarantees that the correct computation was performed. Secure aggregation protocols could be used with VerifiableFL to provide both privacy and integrity guarantees.
VerifiableFL is compatible with a range of FL-specific techniques: robust aggregation (pillutla2022robust) make the FL training process more resilient against integrity attacks; Tolpegin et al. (tolpegin2020data) detect data poisoning attacks using outlier detection. To protect FL training against model poisoning attacks, Krum (blanchard2017machine) and Trimmed-mean (yin2018byzantine) use Byzantine-robust aggregation protocols for secure aggregation. While such protocols require at least a majority of trusted data providers, VerifiableFL’s verifiable claims hold with any number of malicious participants.
Trusted execution for FL. Mo et al. (mo_ppfl_2021) assume that both the FL server and clients are honest-but-curious. They use TEEs to prevent participants from observing the model data during training, thus preventing inference or data reconstruction attacks. Messaoud et al. (valerio21gradsec) address FL client-side privacy threats by protecting gradient updates inside TEEs. Since a subset of the model layers is prevented from gradient leakage, this mitigates against inference attacks. Both of the approaches assume a weaker threat model, without VerifiableFL’s detection of integrity attacks.
CrowdGuard (rieger_close_2022) uses TEEs to achieve integrity-preserving FL. At each training step, data providers use TEEs to evaluate the locally-trained models of every other data provider on their own data and vote regarding poisoned model. Unlike VerifiableFL, CrowdGuard assumes that TEEs offer confidentiality, which makes it vulnerable to side-channel attacks.
Ensuring integrity. Proof-carrying code (necula1997proof) and model checking (clarke1997model) can show that code complies with integrity policies by verifying it against a specification. Since VerifiableFL relies on execution integrity, it can be combined with such approaches.
Other approaches enforce control-flow integrity during execution: C-FLAT (abera2016cflat) and LO-FAT (lo-fat) attest intra-program control flow to ascertain the current execution state of a single node. In contrast, VerifiableFL attests task-level data transformations across distributed participants, allowing for verifiable claims about a final trained FL model. While there are similarities (e.g., use of hashes), the attestation objective is different: in particular, two FL training runs, despite identical control flow attestations, would still produce different models if the datasets are different.
@@ 9. Conclusions
We described VerifiableFL, a new system for executing FL jobs, while enabling an auditor to verify claims about the final trained FL model. VerifiableFL introduces exclaves as a new abstraction for the execution of FL tasks with integrity only, which attest all performed data transformations at runtime. Since executed FL tasks are attested through exclave data records (EDRs), it becomes impossible for participants to deviate from the correct FL training protocol, thus violating model claims, without detection. Exclaves produce EDRs, which can be combined to create an exclave dataflow graph (EDG) that spans the entire FL job. An auditor uses the EDG to verify end-to-end claims about the FL job. We showed that VerifiableFL can be integrated with a state-of-the-art FL framework to detect a wide range of deviations with low runtime overhead.